# thanks: Equal contributorsthanks: Equal contributors

### v2

@@
Fig. 4A, in particular, showcases a suite of designs corresponding to axial shape morphing with $\lambda_{1}$ and $\lambda_{2}$ varied uniformly from $0.8$ (contraction) to $1.3$ (expansion) in a design matrix. The coloring scheme reflects whether the optimized design achieves the target $0.003$ energy barrier. As the coloring indicates, the richness of the design space depends significantly on the shape-morphing. Non-auxetic designs where one side contracts and the other expands appear to be much more amenable to the large energy barriers than auxetic ones. In fact, the most extreme auxetic design – the purely dilation one in the lower right corner of the figure – is only a slight modification of the purely mechanistic rotating squares pattern, even though we optimize for a high energy barrier. This observation suggests that the rotating squares pattern is perhaps the singular template for extreme dilation in quad kirigami. Another interesting point concerns symmetry and non-uniqueness. As illustrated in the design matrix, every optimized design in the upper right quadrant $(\lambda_{1},\lambda_{2})=(x,y)$ is related to one in the lower left $(\lambda_{1},\lambda_{2})=(y,x)$ by a $90^{o}$ rotation. In some of the less extreme cases, however, more than one design achieves the target energy and shape change. We illustrate this point by highlighting a $(\lambda_{1},\lambda_{2})=(0.9,1.2)$ optimized design that distinct from the $(\lambda_{1},\lambda_{2})=(1.2,0.9)$ case shown.
@@
Fig. 4B also showcases a suite of designs, this time for the shear case in Eq. [6], with $\gamma$ evolving uniformly from $0.2$ to $0.95$. Again the design space shrinks as the shear becomes more extreme, making it harder for the optimized design to achieve the target energy barrier $E_{b}^{\text{targ}}=0.003$. Note that the maximum shear in this setting is $\gamma_{\text{max}}=1$, since this shape change takes an effectively square reference domain to a line. Evolving the shear monotonically to this maximum leads to panels that degenerate to lines and slits to parallelograms. A curious yet persistent observation in Fig. 4A-B (and SI Appendix, Section S2.F) is that extremal shape change seems always to correspond to designs with parallelogram slits, i.e., designs known to always posses a single DOF mechanism [54]. Whether this observation suggests a universal relationship between mechanism-based designs and bistable ones remains to be seen, but it is nonetheless compelling evidence of some sort of connection.
@@
Returning to prototyping, Fig. 4C-E highlights the design and fabrication of three examples of extreme shape change: a shear case, a non-auxetic case, and an auxetic, purely dilational case. In all cases, we employ the third fabrication strategy discussed above with pin-joints and laser cut PETG panels, since specimens made this way are easy to produce and exhibit a ”clean” performance. As shown in the figure, each sample is bistable and displays the predicted shape change. The supplementary videos listed and described in SI Appendix Section S6 provide further illustrations of the bistability of these samples.
@@
We end this section by showcasing designs tuned to achieve multiple objectives at the same time. Going back to the prototypical square-to-rectangular transformation studied previously, Fig. 4F shows designs that have been optimized for both a target energy barrier and a target stiffness in its first stable state. Specifically, we fix $c_{b}=1$, $c_{1}=0.002$ and $E_{b}^{\text{targ}}=0.001$ and vary the target stiffness uniformly from $k_{1}^{\text{targ}}=0.15$ to $0.50$ to produce eight optimized designs. As the plot indicates, we have control of both the stiffness and energy barrier over a wide range of the parameter space (from $k_{1}^{\text{targ}}=0.1$ to $0.45$). However, once $k_{1}^{\text{targ}}$ is sufficiently large, the energy barrier can no longer be held fixed. Instead it tilts up, reflecting a tradeoff between high stiffness and low energy.
@@
Figure 5: Heterogeneous designs for complex shape change. (A-C) Example of a pattern designed to morph from a square to a bowtie shape. (A) Mesh grids of the two desired stable states, which give information on the lattice vectors to be prescribed to each unit cell. (B) Optimized pattern in its first and second stable states. (C) Numerical validation of bistability via a geometrically-nonlinear truss model, showing a force-displacement curve and three snapshots of the deformation. Inset (C-1) indicates the boundary conditions for the simulation. (D) Beating heart pattern, in its first and second stable states. (E) Structure designed to morph from a square to a disc, in its first and second stable states.
@@
In the transformation in Fig. 5A, the quad cells of the square mesh are mapped one-to-one to the cells of the bowtie mesh through the lattice transformations $\bm{\ell}_{1}^{R}(i,j)\mapsto\bm{\ell}_{1}^{D}(i,j)$ and $\bm{\ell}_{2}^{R}(i,j)\mapsto\bm{\ell}_{2}^{D}(i,j)$, indexed by $(i,j)$ as shown. While the bowtie mesh is not homogeneous, its lattice vectors vary slowly from cell to cell to produce the overall shape. We exploit these slow variations, and the key fact that the lattice vectors are designer inputs for bistability in Eq. [1], to obtain a kirigami pattern capable of transforming from the square to the bowtie shape at very little stress.
The general idea is as follows (see SI Appendix, Section S3.A for more details). After an initialization step to seed the design of a single kirigami cell, we produce a global kirigami pattern through a marching procedure that is fundamentally local. At each $(i,j)$, indicating corresponding quads in the two meshes, we prescribe the lattice vectors for a bistable design in Eq. [1] as $\bm{\ell}=\bm{\ell}(i,j)$. Next, we choose the remaining DOFs in this equation as a set of minimizers $\mathbf{s}_{1}=\mathbf{s}_{1}(i,j),\mathbf{s}_{2}=\mathbf{s}_{2}(i,j),\bm{\phi}=\bm{\phi}(i,j)$ to Eq. [2] with
@@
where $(\mathbf{s}^{\text{prev}}_{1},\mathbf{s}^{\text{prev}}_{2},\bm{\phi}^{\text{prev}})$ is from a previously computed neighboring cell. Finally, we choose $\mathbf{s}_{3}=\mathbf{s}_{3}(i,j),\ldots,\mathbf{v}=\mathbf{v}(i,j)$ to solve Eq. [1] for the above $(i,j)$ design variables. This recipe furnishes a bistable kirigami cell that takes the shape of the two $(i,j)$-quads as its stable states. Iterating on it produces two global patterns, one with the desired overall square shape and another with the desired bowtie. However, each has small gaps between neighboring unit cells due to the spatial variations of the lattice vectors (see Fig. S4 in the SI Appendix). A final averaging step glues each pattern together and furnishes the kirigami designs for the two shapes shown in Fig. 5B.
The averaging part of the design procedure yields panels in the bowtie that are distorted slightly from their counterparts in the square pattern, meaning that the transformation is not stress-free. To verify bistability, we supplement the procedure with a truss based model of the pattern under loads in Fig. 5C, based on the bar-hinge model of Ref. [52]. The model assumes that the square pattern is the stress-free reference configuration, that bars deform only axially, and that their material is linear elastic (see SI Appendix, Section S7.D). The pattern is then supported by rollers on its left boundary and loaded by a uniform set of horizontal nodal forces $F$ on its right (Fig. 5C-1). The overall horizontal displacement, denoted $d$, increases smoothly under force control until the configuration in Fig 5C-2, where it jumps from $d/|\bm{\ell}_{1}^{R}|\approx 1$ to $d/|\bm{\ell}_{1}^{R}|\approx 1.6$ on a further increase of load. Unloading after the jump, the curve crosses the zero force axis away from the origin, providing a demonstration of bistability. The bowtie shape of Fig. 5C-3, showing small residual strains, is the second equilibrium configuration.
@@
For the heart, we mesh the compact state (see Fig. S5A in the SI Appendix) using the ‘Quasi-Structured Quad’ setting from the freely available software Gmsh [55], and dilate this mesh to obtain the enlarged meshed state. The design procedure iterates through the cells in these meshes to produce the kirigami pattern. We have studied a variety of dilation factors $\lambda\geq 1$. Increasing this factor monotonically leads to a compact state where the slits degenerate to lines (closed slits), and a dilated state where the slits tend towards being fully open. At sufficiently large values for $\lambda$, the panels begin to overlap in the compact state, violating the inequality constraints in Eq. [2]. Fig. 5D shows a $\lambda=1.25$ design obtained by our methods. This level of dilation is actually quite impressive given the heterogeneity of the cells and certain basic limitations of quad kirigami. The famed rotating squares pattern, for instance, transforms from its fully closed to fully open state by a uniform dilation of $\lambda_{rs}=\sqrt{2}\approx 1.41$, which likely sets the theoretical upper bound on dilation for these types of patterns.
Our last demonstration of heterogeneous morphing in Fig. 5E showcases a square-to-disc kirigami pattern. The quad mesh precursors are obtained by discretizing the square domain uniformly on $(-1,1)^{2}$ using a $40\times 40$ set of square cells, and then deforming this mesh smoothly to the unit disc by the Elliptical Grid mapping $(x,y)\mapsto(x\sqrt{1-y^{2}/2},y\sqrt{1-x^{2}/2})$. This mapping spreads the mesh distortions smoothly from the interior to four singular points, corresponding to the deformed corners of the square (see Fig. S5A in the SI Appendix). The design procedure accounts for this spreading by producing cells in the square whose slit area increases dramatically from its center to its corners; the dichotomy is reversed in the disc state. As shown in Fig. 5E, the corners of the square essentially collapse inward, closing their slits, to achieve the disc as the second stable state. A key ingredient to this example’s impressive display of shape change is the scalability of our methods. Since the design procedure is completely local, we can produce kirigami designs with target shapes that involve thousands of unit cells in a matter of minutes on a standard laptop.
@@ Conclusions
In summary, we have introduced a set of principles to rationally design and optimize bistable planar kirigami metamaterials, starting from a well known, non-bistable template. Importantly, the optmization tools incorporate a reduced-order elastic model, enabling on-demand design guidance for bistable kirigami with a rich variety of shape morphing capabilities and energy landscapes – it provides experimentalists access to new designs in a matter of minutes. Optimized bistable designs are exemplified through a variety of fabrication strategies, with experimental results that largely match the shape change of the theory and the trends in the target elastic properties. Open questions remain. The fine details of fabrication introduce features not present in the current design framework, e.g., hinge elasticity, beam buckling or panel bending. One challenge is to adapt the framework to account for these details, while still remaining an efficient design tool. Another avenue concerns generality. The basic ingredients to our framework are geometric perturbations of a unit cell and intercell compatibility constraints. As these ingredients are found across metamaterial templates, design formulas that hard encode bistability, like Eq. [1], are perhaps ripe for discovery in a wide range of metamaterials.
@@ Acknowledgments
The panel loop conditions for this state are trivially implied by their counterparts in the second line of in Eq. [S2] and are thus not a constraint. Periodicity implies its own set of equality constraints on these parameters. We deduce from Fig. S1A and B that reference Bravais lattice vectors satisfy
@@
Eqs. [S2-S5] complete the description of the equality constraints governing a compatible design for bistable kirigami.
@@
The last set of convexity conditions are for the four slits in the second stable state. These conditions are given by replacing the $\mathbf{s}_{1},\ldots,\mathbf{v}_{4}$ vectors in Eq. [S6] with their deformed counterparts, i.e.,
@@
Similar to the loop conditions, the convexity conditions of the deformed panels are implied by the reference versions in Eq. [S7]. While there are most certainly redundancies in the 48 inequalities in Eqs. [S6-S8], this issue is not so important as to be worth refining the equations.
@@
Eqs. [S6-S9] enumerate all the inequality constraints associated to a compatible design. Overall, they form a list of 52 nonlinear inequalities of the form
@@
for organizational purposes. This list of inequalities, combined with Eqs. [S2-S5], furnish the necessary and sufficient conditions for a cell-based quad kirigami metamaterial with at least two stable states of specified lattice vectors $\bm{\ell}_{1}^{R},\bm{\ell}_{2}^{R},\bm{\ell}_{1}^{D},\bm{\ell}_{2}^{D}$.
@@
We now derive the key design formula parameterizing the equality constraints in Eqs. [S2-S5]. First eliminate $\mathbf{v}_{1},\ldots,\mathbf{v}_{4}$ via the parameterizations
@@
Solving Eqs. [S15] and [S16] completes the description.
@@
Direct substitution of Eq. [S17] into Eq. [S16] leads to the conditions
@@
where the second equality in each formula is the result of a careful manipulation using the correspondence between design vectors $\mathbf{s}_{1},\ldots,\mathbf{v}_{4}$ and points $\mathbf{x}_{ij}$ in Fig. S1A and Fig. S2A, as well as the loop conditions in Eq. [S2].
@@
The loop conditions in Eqs. [S2] imply that, on Taylor expansion, the first stable state satisfies

### v3

@@ Abstract
Fig. 4A, in particular, showcases a suite of designs corresponding to axial shape morphing with $\lambda_{1}$ and $\lambda_{2}$ varied uniformly from $0.8$ (contraction) to $1.3$ (expansion) in a design matrix. The coloring scheme reflects whether the optimized design achieves the target $0.003$ energy barrier. As the coloring indicates, the richness of the design space depends significantly on the shape-morphing. Non-auxetic designs where one side contracts and the other expands appear to be much more amenable to the large energy barriers than auxetic ones. In fact, the most extreme auxetic design – the purely dilation one in the lower right corner of the figure – is only a slight modification of the purely mechanistic rotating squares pattern, even though we optimize for a high energy barrier. This observation suggests that the rotating squares pattern is perhaps the singular template for extreme dilation in quad kirigami. Another interesting point concerns symmetry and non-uniqueness. As illustrated in the design matrix, every optimized design in the upper right quadrant $(\lambda_{1},\lambda_{2})=(x,y)$ is related to one in the lower left $(\lambda_{1},\lambda_{2})=(y,x)$ by a $90^{o}$ rotation. These rotated designs are shaded in Fig. 4A. In some of the less extreme cases, however, more than one design achieves the target energy and shape change. We illustrate this point by highlighting a $(\lambda_{1},\lambda_{2})=(0.9,1.2)$ optimized design that is distinct from the $(\lambda_{1},\lambda_{2})=(1.2,0.9)$ case shown.
@@
Fig. 4B also showcases a suite of designs, this time for the shear case in Eq. [6], with $\gamma$ evolving uniformly from $0.2$ to $0.95$. Again the design space shrinks as the shear becomes more extreme, making it harder for the optimized design to achieve the target energy barrier $E_{b}^{\text{targ}}=0.003$. Note that the maximum shear in this setting is $\gamma_{\text{max}}=1$, since this shape change takes an effectively square reference domain to a line. Evolving the shear monotonically to this maximum leads to panels that degenerate to lines and slits to parallelograms. A curious yet persistent observation in Fig. 4A-B (and Supplemental Material, Section S2.F[56]) is that extremal shape change seems always to correspond to designs with parallelogram slits, i.e., designs known to always posses a single DOF mechanism [53]. Whether this observation suggests a universal relationship between mechanism-based designs and bistable ones remains to be seen, but it is nonetheless compelling evidence of some sort of connection.
@@
Returning to prototyping, Fig. 4C-E highlights the design and fabrication of three examples of extreme shape change: a shear case, a non-auxetic case, and an auxetic, purely dilational case. In all cases, we employ the third fabrication strategy discussed above with pin-joints and laser cut PETG panels, since specimens made this way are easy to produce and exhibit a ”clean” performance. As shown in the figure, each sample is bistable and displays the predicted shape change. The supplementary videos listed and described in Supplemental Material, Section S6[56] provide further illustrations of the bistability of these samples.
Figure 5: Additional design explorations. Theoretical energy landscapes of patterns designed to display various target stiffness $k_{1}^{\text{targ}}$, while maintaining a barrier around 0.001. In the legend, we show the first stable states of all such patterns.
@@
We end this section by showcasing designs tuned to achieve multiple objectives at the same time. Going back to the prototypical square-to-rectangular transformation studied previously, Fig. 5 shows designs that have been optimized for both a target energy barrier and a target stiffness in its first stable state. Specifically, we fix $c_{b}=1$, $c_{1}=0.002$ and $E_{b}^{\text{targ}}=0.001$ and vary the target stiffness uniformly from $k_{1}^{\text{targ}}=0.15$ to $0.50$ to produce eight optimized designs. As the plot indicates, we have control of both the stiffness and energy barrier over a wide range of the parameter space (from $k_{1}^{\text{targ}}=0.1$ to $0.45$). However, once $k_{1}^{\text{targ}}$ is sufficiently large, the energy barrier can no longer be held fixed. Instead it tilts up, reflecting a tradeoff between high stiffness and low energy.
@@
Figure 6: Heterogeneous designs for complex shape change. (A-C) Example of a pattern designed to morph from a square to a bowtie shape. (A) Mesh grids of the two desired stable states, which give information on the lattice vectors to be prescribed to each unit cell. (B) Optimized pattern in its first and second stable states. (C) Numerical validation of bistability via a geometrically-nonlinear truss model, showing a force-displacement curve with arrows indicating the direction of loading/unloading and three snapshots of the deformation. Inset (C-1) indicates the boundary conditions for the simulation. (D) Beating heart pattern, in its first and second stable states. (E) Structure designed to morph from a square to a disc, in its first and second stable states.
@@
In the transformation in Fig. 6A, the quad cells of the square mesh are mapped one-to-one to the cells of the bowtie mesh through the lattice transformations $\bm{\ell}_{1}^{R}(i,j)\mapsto\bm{\ell}_{1}^{D}(i,j)$ and $\bm{\ell}_{2}^{R}(i,j)\mapsto\bm{\ell}_{2}^{D}(i,j)$, indexed by $(i,j)$ as shown. While the bowtie mesh is not homogeneous, its lattice vectors vary slowly from cell to cell to produce the overall shape. We exploit these slow variations, and the key fact that the lattice vectors are designer inputs for bistability in Eq. [LABEL:eq:theorem1], to obtain a kirigami pattern capable of transforming from the square to the bowtie shape at very little stress.
The general idea is as follows (see Supplemental Material, Section S3[56] for more details). After an initialization step to seed the design of a single kirigami cell, we produce a global kirigami pattern through a marching procedure that is fundamentally local. At each $(i,j)$, indicating corresponding quads in the two meshes, we prescribe the lattice vectors for a bistable design in Eq. [LABEL:eq:theorem1] as $\bm{\ell}=\bm{\ell}(i,j)$. Next, we choose the remaining DOFs in this equation as a set of minimizers $\mathbf{s}_{1}=\mathbf{s}_{1}(i,j),\mathbf{s}_{2}=\mathbf{s}_{2}(i,j),\bm{\phi}=\bm{\phi}(i,j)$ to Eq. [2] with
@@
where $(\mathbf{s}^{\text{prev}}_{1},\mathbf{s}^{\text{prev}}_{2},\bm{\phi}^{\text{prev}})$ is from a previously computed neighboring cell. Finally, we choose $\mathbf{s}_{3}=\mathbf{s}_{3}(i,j),\ldots,\mathbf{v}=\mathbf{v}(i,j)$ to solve Eq. [LABEL:eq:theorem1] for the above $(i,j)$ design variables. This recipe furnishes a bistable kirigami cell that takes the shape of the two $(i,j)$-quads as its stable states. Iterating on it produces two global patterns, one with the desired overall square shape and another with the desired bowtie. However, each has small gaps between neighboring unit cells due to the spatial variations of the lattice vectors (see Fig. S5 in the Supplemental Material[56]). A final averaging step glues each pattern together and furnishes the kirigami designs for the two shapes shown in Fig. 6B.
The averaging part of the design procedure yields panels in the bowtie that are distorted slightly from their counterparts in the square pattern, meaning that the transformation is not stress-free. To verify bistability, we supplement the procedure with a truss based model of the pattern under loads in Fig. 6C, based on the bar-hinge model of Ref. [57]. The model assumes that the square pattern is the stress-free reference configuration, that bars deform only axially, and that their material is linear elastic (see Supplemental Material, Section S7.D[56]). The pattern is then supported by rollers on its left boundary and loaded by a uniform set of horizontal nodal forces $F$ on its right (Fig. 6C-1). The overall horizontal displacement, denoted $d$, increases smoothly under force control until the configuration in Fig 6C-2, where it jumps from $d/|\bm{\ell}_{1}^{R}|\approx 1$ to $d/|\bm{\ell}_{1}^{R}|\approx 1.6$ on a further increase of load. Unloading after the jump, the curve crosses the zero force axis away from the origin, providing a demonstration of bistability. The bowtie shape of Fig. 6C-3, showing small residual strains, is the second equilibrium configuration.
@@
For the heart, we mesh the compact state (see Fig. S7A in the Supplemental Material[56]) using the ‘Quasi-Structured Quad’ setting from the freely available software Gmsh [59], and dilate this mesh to obtain the enlarged meshed state. The design procedure iterates through the cells in these meshes to produce the kirigami pattern. We have studied a variety of dilation factors $\lambda\geq 1$. Increasing this factor monotonically leads to a compact state where the slits degenerate to lines (closed slits), and a dilated state where the slits tend towards being fully open. At sufficiently large values for $\lambda$, the panels begin to overlap in the compact state, violating the inequality constraints in Eq. [2]. Fig. 6D shows a $\lambda=1.25$ design obtained by our methods. This level of dilation is actually quite impressive given the heterogeneity of the cells and certain basic limitations of quad kirigami. The rotating squares pattern, for instance, transforms from its fully closed to fully open state by a uniform dilation of $\lambda_{rs}=\sqrt{2}\approx 1.41$, which likely sets the theoretical upper bound on dilation for these types of patterns.
Our last demonstration of heterogeneous morphing in Fig. 6E showcases a square-to-disc kirigami pattern. The quad mesh precursors are obtained by discretizing the square domain uniformly on $(-1,1)^{2}$ using a $40\times 40$ set of square cells, and then deforming this mesh smoothly to the unit disc by the Elliptical Grid mapping $(x,y)\mapsto(x\sqrt{1-y^{2}/2},y\sqrt{1-x^{2}/2})$. This mapping spreads the mesh distortions smoothly from the interior to four singular points, corresponding to the deformed corners of the square (see Fig. S7B in the Supplemental Material[56]). The design procedure accounts for this spreading by producing cells in the square whose slit area increases dramatically from its center to its corners; the dichotomy is reversed in the disc state. As shown in Fig. 6E, the corners of the square essentially collapse inward, closing their slits, to achieve the disc as the second stable state. A key ingredient to this example’s impressive display of shape change is the scalability of our methods. Since the design procedure is completely local, we can produce kirigami designs with target shapes that involve thousands of unit cells in a matter of minutes on a standard laptop.
@@ Conclusions
In summary, we have introduced a set of principles to rationally design and optimize bistable planar kirigami metamaterials, starting from a well known, non-bistable template. Importantly, the optmization tools incorporate a reduced-order elastic model, enabling on-demand design guidance for bistable kirigami with a rich variety of shape morphing capabilities and energy landscapes – it provides experimentalists access to new designs in a matter of minutes. Optimized bistable designs are exemplified through a variety of fabrication strategies, with experimental results that largely match the shape change of the theory and the trends in the target elastic properties. Open questions remain. The fine details of fabrication introduce features not present in the current design framework, e.g., hinge elasticity, beam buckling or panel bending. One challenge is to adapt the framework to account for these details, while still remaining an efficient design tool. Another avenue concerns generality. The basic ingredients to our framework are geometric perturbations of a unit cell and intercell compatibility constraints. As these ingredients are found across metamaterial templates, design formulas that hard encode bistability, like Eq. [LABEL:eq:theorem1], are perhaps ripe for discovery in a wide range of metamaterials.
@@ Acknowledgments
The panel loop conditions for this state are trivially implied by their counterparts in the second line of in Eq. [LABEL:eq:LoopRef] and are thus not a constraint. Periodicity implies its own set of equality constraints on these parameters. We deduce from Fig. S1A and B that reference Bravais lattice vectors satisfy
@@
Eqs. [LABEL:eq:LoopRef-S5] complete the description of the equality constraints governing a compatible design for bistable kirigami.
@@
The last set of convexity conditions are for the four slits in the second stable state. These conditions are given by replacing the $\mathbf{s}_{1},\ldots,\mathbf{v}_{4}$ vectors in Eq. [LABEL:eq:slitRefConvexity] with their deformed counterparts, i.e.,
@@
Similar to the loop conditions, the convexity conditions of the deformed panels are implied by the reference versions in Eq. [LABEL:eq:panelRefConvexity]. While there are most certainly redundancies in the 48 inequalities in Eqs. [LABEL:eq:slitRefConvexity-LABEL:eq:slitDefConvexity], this issue is not so important as to be worth refining the equations.
@@
Eqs. [LABEL:eq:slitRefConvexity-S9] enumerate all the inequality constraints associated to a compatible design. Overall, they form a list of 52 nonlinear inequalities of the form
@@
for organizational purposes. This list of inequalities, combined with Eqs. [LABEL:eq:LoopRef-S5], furnish the necessary and sufficient conditions for a cell-based quad kirigami metamaterial with at least two stable states of specified lattice vectors $\bm{\ell}_{1}^{R},\bm{\ell}_{2}^{R},\bm{\ell}_{1}^{D},\bm{\ell}_{2}^{D}$.
@@
We now derive the key design formula parameterizing the equality constraints in Eqs. [LABEL:eq:LoopRef-S5]. First eliminate $\mathbf{v}_{1},\ldots,\mathbf{v}_{4}$ via the parameterizations
@@
Solving Eqs. [S15] and [LABEL:eq:AfterManip2] completes the description.
@@
Direct substitution of Eq. [S17] into Eq. [LABEL:eq:AfterManip2] leads to the conditions
@@
where the second equality in each formula is the result of a careful manipulation using the correspondence between design vectors $\mathbf{s}_{1},\ldots,\mathbf{v}_{4}$ and points $\mathbf{x}_{ij}$ in Fig. S1A and Fig. S2A, as well as the loop conditions in Eq. [LABEL:eq:LoopRef].
@@
The loop conditions in Eqs. [LABEL:eq:LoopRef] imply that, on Taylor expansion, the first stable state satisfies